### v1

@@ 6 Systematic Experiments
Here, we conduct systematic experiments to demonstrate the superiority of the proposed algorithm using the Zero rule over algorithms in Avis and Fukuda [4] and Moss [25]. We independently implement the algorithms in Avis and Fukuda [4] and Moss [25] by following their algorithmic structures. We implement our algorithm according to the outline shown in Algorithm 6. We generate four different arrangements: unit hypercubes, unit cubes with a cone being cut i.e., adding the equation $y_{1}+y_{2}+\cdots+y_{d}=1.5$ with the unit cube, random hyperplanes defined by random integers as coefficients, and a three-layer ReLU neural network. The merit of the former is the ground truth is known, which can assist us to examine if the compared algorithms can enumerate all vertices upon completion of the algorithm. The latter is to show the comparison in the general case. All our examples and Matlab code are publicly available [^1] https://github.com/Github-DongZelin/Examples-and-Command-of-the-algorithm for readers’ free download and use.
Unit hypercubes: Table 2 presents different algorithms’ performance over unit hypercubes ranging from $\mathbb{R}^{4}$ to $\mathbb{R}^{10}$. It is surprising that Avis and Fukuda [4]’s algorithm cannot find all vertices in such a direct case. The reason is that this algorithm skips many vertices and stops early. Specifically, in certain steps, applying the Criss-Cross rule jumps into the initial dictionary and selects much later entries to pivot, which makes the algorithm terminate earlier without finding all vertices.
$\mathbb{R}^{4}$
$\mathbb{R}^{5}$
$\mathbb{R}^{6}$
$\mathbb{R}^{7}$
$\mathbb{R}^{8}$
$\mathbb{R}^{9}$
$\mathbb{R}^{10}$
Truth
16
32
64
128
256
512
1024
@@
16
32
64
128
256
512
1024
Avis and Fukuda [4]
5
6
7
8
9
10
11
Ours
16
32
64
128
256
512
1024
Table 2: The number of vertices found by different algorithms over unit hypercubes from $\mathbb{R}^{4}$ to $\mathbb{R}^{10}$.
Unit hypercubes with a cone being cut: Table 3 presents different algorithms’ performance over unit hypercubes with a cone being cut ranging from $\mathbb{R}^{4}$ to $\mathbb{R}^{10}$. In these examples, Avis and Fukuda’s algorithm [4] once again fails to find all vertices. Unlike the previous performance, their algorithm goes into a loop, causing it to fail in finding all vertices. After the table, Figures 12 and Figure 12 demonstrate the process of different algorithms in finding all vertices in $\mathbb{R}^{9}$ and $\mathbb{R}^{10}$, respectively; our method is better.
$\mathbb{R}^{4}$
$\mathbb{R}^{5}$
$\mathbb{R}^{6}$
$\mathbb{R}^{7}$
$\mathbb{R}^{8}$
$\mathbb{R}^{9}$
$\mathbb{R}^{10}$
@@
48
112
256
576
1280
2816
6144
Avis and Fukuda [4]
6
7
8
9
10
11
12
Ours
48
112
256
576
1280
2816
6144
Table 3: The number of vertices found by different algorithms over unit hypercubes with a cone being cut from $\mathbb{R}^{4}$ to $\mathbb{R}^{10}$.
Figure 11: cube with a cone being cut in $\mathbb{R}^{9}$
Figure 11: cube with a cone being cut in $\mathbb{R}^{9}$
Figure 12: cube with a cone being cut in $\mathbb{R}^{10}$
Random arrangement. By taking random integers as coefficients of arrangements, a total of 25 cases are generated by varying from $\mathbb{R}^{2}$ to $\mathbb{R}^{6}$ and from 7 to 15 hyperplanes. Each case has five runs to ensure the reliability of the results. Table 4 displays the average number of vertices found by three algorithms. Table 5 shows the average computation time. For each case, we set 5000 seconds as the critical value. If the time in a case exceeds this value, we denote it as $>5000$. The algorithm falls into a cycle in processing all examples when using Avis and Fukuda’s method.
$d\backslash n$
Method
7
9
11
13
15
$\mathbb{R}^{2}$
@@
21
35
53.8
75.2
99.6
Avis and Fukuda [4]
8.2
7.4
8.6
7
8
Ours
21
35
53.8
75.2
99.6
$\mathbb{R}^{3}$
@@
35
83.8
163
283.8
445.2
Avis and Fukuda [4]
9
8.2
8.8
16.4
11.2
Ours
35
83.8
163
283.8
445.2
$\mathbb{R}^{4}$
@@
35
126
330
714
1364.4
Avis and Fukuda [4]
9
9.8
12.4
12
11
Ours
35
126
330
714
1364.4
$\mathbb{R}^{5}$
@@
21
126
462
1287
3002.6
Avis and Fukuda [4]
7.6
8.6
19.2
10.4
14
Ours
21
126
462
1287
3002.6
$\mathbb{R}^{6}$
@@
$d\backslash n$
Method
7
9
11
13
15
$\mathbb{R}^{2}$
@@
1.8357
3.4629
6.3672
10.339
15.709
Avis and Fukuda [4]
$>5000$
$>5000$
$>5000$
$>5000$
$>5000$
Ours
1.7081
3.5555
6.5597
10.531
16.288
$\mathbb{R}^{3}$
@@
2.8036
8.2725
19.666
39.980
61.593
Avis and Fukuda [4]
$>5000$
$>5000$
$>5000$
$>5000$
$>5000$
Ours
2.6670
8.0951
19.207
39.196
70.270
$\mathbb{R}^{4}$
@@
2.6872
12.382
40.707
102.05
229.00
Avis and Fukuda [4]
$>5000$
$>5000$
$>5000$
$>5000$
$>5000$
Ours
22.5885
12.051
38.789
100.72
217.85
$\mathbb{R}^{5}$
Moss [25]
1.4985
12.137
55.207
186.65
502.96
Avis and Fukuda [4]
$>5000$
$>5000$
$>5000$
$>5000$
$>5000$
Ours
1.4672
12.059
54.835
184.75
483.07
$\mathbb{R}^{6}$
Moss [25]
0.4936
7.6735
54.772
246.42
960.63
Avis and Fukuda [4]
$>5000$
$>5000$
$>5000$
$>5000$
$>5000$
Ours
0.4702
7.9303
54.968
246.02
817.43
@@
Figure 13: average number of vertex $\mathbb{R}^{5}$
Figure 13: average number of vertex $\mathbb{R}^{5}$
Figure 14: average number of vertex $\mathbb{R}^{6}$
Using the Zero rule significantly improves Moss’ method, while reversing the Criss-Cross rule did not finish all cases well. Additionally, from each column in Table 5, Figure 14 and Figure 14, it can be observed that as the number of dimensions and vertices increases, the improvement becomes even more pronounced for any number of hyperplanes. This aligns with our complexity analysis, indicating that if the number of hyperplanes and the number of dimensions continuously increase, the Zero rule can save more computational overhead.
A three-layer ReLU neural network: We have built a neural network model with an input layer consisting of $7$ nodes to handle data from $\mathbb{R}^{5}$. Each node in the input layer will receive a $5$-dimensional input. The hidden layer comprises $5$ nodes, each utilizing a weight vector with $7$ dimensions along with a bias term. These layers employ the Rectified Linear Unit (ReLU) activation function to introduce non-linearity into the network. In this network, all weight vectors and bias terms are randomly generated using the rand function in MATLAB. Firstly, we use auxiliary linear programming to find the partition of the seven neurons in the input layer, resulting in a total of $119$ cells in $\mathbb{R}^{5}$. As the activation state of each cell is different, $5$ hyperplanes corresponding to the nodes in the hidden layer will also be different. Next, we use algorithms to find out the total number of vertex of these $119$ different hyperplane arrangements with $12=7+5$ hyperplanes in $\mathbb{R}^{5}$, meanwhile, during the execution of each algorithm, we also select and count the vertices on the cell that corresponding to the activation states and the time taken for these computations. The naming order of the cells is only related to the sequence in which we found them, and partly of the results are provided in Tables 6 and 7 (Moss: Moss [25]; AF: Avis and Fukuda [4]) and the overall result can be found in our link [^2] https://github.com/Github-DongZelin/Examples-and-Command-of-the-algorithm, $-1$ in the first cell means none of the seven nodes in the input layer is active.
Method 1 4 7
feasible
all
feasible
all
feasible
all
0
[25]
$-1$
$-1$
7
246
4
246
[4]
$-1$
$-1$
1
12
1
17
Ours
$-1$
$-1$
7
246
4
246
10
[25]
9
246
2
456
7
96
[4]
1
7
1
8
1
7
Ours
9
246
2
456
7
96
20
[25]
5
246
6
666
10
456
[4]
1
9
1
16
1
6
Ours
5
246
6
666
10
456
30
[25]
8
666
3
246
1
666
[4]
1
11
2
16
1
7
Ours
8
666
3
246
1
666
40
[25]
6
666
3
666
4
456
AF
2
3
2
11
2
8
Ours
6
666
3
666
4
456
50
[25]
2
792
5
666
4
792
[4]
1
17
2
6
1
13
Ours
2
792
5
666
4
792
60
[25]
6
246
3
246
4
666
[4]
3
14
1
7
2
8
Ours
6
246
3
246
4
666
70
[25]
9
666
4
792
9
456
[4]
3
18
2
7
5
21
Ours
9
666
4
792
9
456
80
[25]
7
666
10
666
7
666
[4]
3
8
2
6
2
7
Ours
7
666
10
666
7
666
90
[25]
4
246
5
666
2
792
[4]
1
11
2
3
1
11
Ours
4
246
5
666
2
792
100
[25]
7
792
3
792
4
666
[4]
1
6
1
6
2
4
Ours
7
792
3
792
4
666
110
[25]
3
792
8
792
5
792
[4]
2
10
2
7
1
22
Ours
3
792
8
792
5
792
Method
1
4
7
11
14
17
0
Moss
$-1$
39.224
36.991
36.848
69.985
13.7739
AF
$>5000$
$>5000$
$>5000$
$>5000$
$>5000$
$>5000$
Ours
$-1$
34.216
37.636
35.147
67.154
14.3105
20
Moss
36.215
103.53
70.071
102.83
36.260
102.93
AF
$>5000$
$>5000$
$>5000$
$>5000$
$>5000$
$>5000$
Ours
36.252
93.205
66.960
95.646
35.053
94.164
40
Moss
104.52
103.48
69.086
127.26
104.40
127.10
AF
$>5000$
$>5000$
$>5000$
$>5000$
$>5000$
$>5000$
Ours
93.487
93.798
67.129
110.63
94.500
112.92
60
Moss
36.556
38.099
104.10
105.85
125.58
70.180
AF
$>5000$
$>5000$
$>5000$
$>5000$
$>5000$
$>5000$
Ours
37.892
36.202
94.394
96.062
111.79
67.700
80
Moss
104.75
104.90
104.78
36.840
104.45
125.81
AF
$>5000$
$>5000$
$>5000$
$>5000$
$>5000$
$>5000$
Ours
95.837
97.517
96.017
36.358
93.960
111.37
100
Moss
125.51
125.95
104.36
126.42
127.87
126.05
AF
$>5000$
$>5000$
$>5000$
$>5000$
$>5000$
$>5000$
Ours
110.67
110.24
95.478
110.19
111.65
109.63
Table 7: The computation time consumed by different algorithms.
From Table 6, it can be seen that the vertices in these hyperplane arrangements are highly degenerate. In a simple arrangement formed by 12 hyperplanes in $\mathbb{R}^{5}$, there should be $492$ vertices. However, only a small portion of these examples reach this value. Therefore, we can also observe the effectiveness of our method in highly degenerate hyperplane distributions from this series of examples. The following Figures 16-22 demonstrate the process of searching vertices in four randomly selected examples, those figures show that when searching vertices of arrangement, our algorithm is better than the random generate integer coefficient arrangement in previous, but when searching vertices on the cell, Their performance is similar and less efficient, but Avis and Fukuda’s method, while not finding all the vertices, is highly efficient in generating vertices on the cell.
Figure 15: vertices in 4-th arrangement
Figure 15: vertices in 4-th arrangement
Figure 16: vertices in 4-th ell
Figure 17: vertices in 47-th arrangement
Figure 18: vertices in 47-st arrangement
Figure 19: vertices in 81-st arrangement
Figure 20: number of vertex in 81-st cell
Figure 21: vertices in 114-th arrangement
Figure 17: vertices in 47-th arrangement
Figure 18: vertices in 47-st arrangement
Figure 19: vertices in 81-st arrangement
Figure 20: number of vertex in 81-st cell
Figure 21: vertices in 114-th arrangement
Figure 22: vertices in 114-th cell
7 Discussion
The above experiments demonstrate that the VE algorithm using the Criss-Cross rule cannot enumerate all vertices. Can we remedy this problem? The reason why using the Criss-Cross rule fails in certain cases is the unrigorous examination of valid reversibility. Here, we complement the Criss-Cross rule by upgrading its if condition to the if-and-only-if condition.
Proposition 7.1.
Let $x_{B}=\bar{A}x_{N}$ be an arbitrary dictionary with the objective function. Then $(s,r)$, where $s\in B_{\neq f}$ and $r\in N_{\neq g}$, is a valid reverse pivot of the Criss-Cross rule if and only if either (a)

### v2

@@ 6 Systematic Experiments
Here, we demonstrate the superiority of our algorithm through systematic experiments. We compare our algorithm with Enhanced AF, AF, and Moss on the following four types of hyperplane arrangements: an arrangement containing $2d$ hyperplanes that enclose a unit cube in $\mathbb{R}^{d}$; $12d+$ 1 hyperplanes that enclose a truncated unit cube in $\mathbb{R}^{d}$, i.e., adding the hyperplane $y_{1}+y_{2}+\cdots+y_{d}=1.5$ with the unit cube; 125 different hyperplane arrangements randomly generated by MATLAB; and a family of arrangements that consist of all possible arrangements generated by a three-layer ReLU neural network with weights randomly generated. For each example, we list the number of vertices found by four algorithms and the run time. We denote the time spent exceeding 10,000 seconds as $>$9,999e purpose of the first two arrangements is that the ground truth is known, which can assist us to examine if the compared algorithms can enumerate all vertices upon completion of the algorithm. The third is to show the comparison in the general case. The last one is to show the utility of the proposed algorithm in artificial networks. All our examples and code are publicly available [^1] https://github.com/Github-DongZelin/Examples-and-Command-of-the-algorithm for readers’ free download and use.
Unit hypercubes: Tables 2 and 3 presents different algorithms’ performance over unit hypercubes ranging from $\mathbb{R}^{2}$ to $\mathbb{R}^{8}$.
Method $\mathbb{R}^{2}$ $\mathbb{R}^{3}$ $\mathbb{R}^{4}$ $\mathbb{R}^{5}$ $\mathbb{R}^{6}$ $\mathbb{R}^{7}$ $\mathbb{R}^{8}$
Truth 4 8 16 32 64 128 256
Ours 4 8 16 32 64 128 256
AF [2]
Enhanced AF
@@
Table 2: The number of vertices found by four algorithms over unit hypercubes from $\mathbb{R}^{2}$ to $\mathbb{R}^{8}$.
Method $\mathbb{R}^{2}$ $\mathbb{R}^{3}$ $\mathbb{R}^{4}$ $\mathbb{R}^{5}$ $\mathbb{R}^{6}$ $\mathbb{R}^{7}$ $\mathbb{R}^{8}$
Ours 0.2201 0.6535 1.9113 4.8336 12.264 30.774 73.478
AF [2] 0.5261 1.5073 4.0578 10.476 26.228 64.643 150.81
Enhanced AF 0.4023 1.0381 2.7246 6.9270 17.287 43.149 105.63
Moss [29] 0.3520 0.8673 2.5098 8.0306 30.945 129.95 628.63
Table 3: The time consuming by four algorithms over unit hypercubes from $\mathbb{R}^{2}$ to $\mathbb{R}^{8}$.
Unit hypercubes with a cone being cut: Tables 5 and 4 presents different algorithms’ performance over this arrangement ranging from $\mathbb{R}^{2}$ to $\mathbb{R}^{8}$.
Method $\mathbb{R}^{2}$ $\mathbb{R}^{3}$ $\mathbb{R}^{4}$ $\mathbb{R}^{5}$ $\mathbb{R}^{6}$ $\mathbb{R}^{7}$ $\mathbb{R}^{8}$
Truth 8 20 48 112 256 576 1280
Ours 8 20 48 112 256 576 1280
AF [2]
Enhanced AF
@@
Table 4: The number of vertices found by four algorithms over unit hypercubes with a cone being cut from $\mathbb{R}^{2}$ to $\mathbb{R}^{8}$.
Method $\mathbb{R}^{2}$ $\mathbb{R}^{3}$ $\mathbb{R}^{4}$ $\mathbb{R}^{5}$ $\mathbb{R}^{6}$ $\mathbb{R}^{7}$ $\mathbb{R}^{8}$
Ours 0.4817 1.5661 5.1182 15.792 46.351 129.48 378.05
AF [2] 1.5312 4.9375 15.173 45.272 124.24 340.59 912.61
Enhanced AF 1.0575 2.8983 8.7185 25.462 71.393 195.76 547.49
Moss [29] 0.9378 3.2863 16.060 96.004 647.23 4314.6 $>$$9$
Table 5: The time consuming by four algorithms over unit hypercubes with a cone being cut from $\mathbb{R}^{2}$ to $\mathbb{R}^{8}$.
Random arrangement. By taking random integers as coefficients of arrangements, a total of 25 cases are generated by varying from $\mathbb{R}^{2}$ to $\mathbb{R}^{6}$ and from 8 to 16 hyperplanes. Each case has five runs to ensure the reliability of the results. Table 6 displays the average number of vertices found by three algorithms. Table 7 shows the average computation time.
$d\backslash n$ Method 8 10 12 14 16
$\mathbb{R}^{2}$ Ours 27.2 42 63.8 88.4 110.6
AF [2]
Enhanced AF
@@
$\mathbb{R}^{3}$ Ours 55.6 119.2 218 362.8 557.6
AF [2]
Enhanced AF
@@
$\mathbb{R}^{4}$ Ours 69.6 210 494.6 997 1818.8
AF [2]
Enhanced AF
@@
$\mathbb{R}^{5}$ Ours 56 252 792 2001 4367.6
AF [2]
Enhanced AF
@@
$\mathbb{R}^{6}$ Ours 28 210 924 3003 8008
AF [2]
Enhanced AF
@@
$d\backslash n$ Method 8 10 12 14 16
$\mathbb{R}^{2}$ Ours 1.5421 2.5362 3.8307 5.4066 7.0743
AF [2] 10.177 20.936 38.954 61.440 87.521
Enhanced AF 4.8326 8.9556 15.564 24.147 33.960
Moss [29] 3.7805 7.6048 14.133 24.475 37.145
$\mathbb{R}^{3}$ Ours 4.268 9.7160 18.370 29.983 47.327
AF [2] 25.643 72.388 171.80 321.27 580.23
Enhanced AF 11.032 28.791 62.738 116.36 202.30
Moss [29] 16.572 65.205 217.43 593.14 1524.9
$\mathbb{R}^{4}$ Ours 6.1188 20.059 50.243 112.76 369.08
AF [2] 32.303 135.14 416.99 1098.1 2533.7
Enhanced AF 13.399 51.728 147.911 378.34 955.68
Moss [29] 34.254 317.89 1960.9 8526.4 $>$$9$
$\mathbb{R}^{5}$ Ours 5.9546 30.227 103.30 401.13 1115.9
AF[2] 28.430 181.00 737.39 2364.4 6653.0
Enhanced AF 12.506 67.570 264.56 843.64 2491.2
Moss [29] 32.466 1387.3 9736.0 $>$$9$ $>$$9$
$\mathbb{R}^{6}$ Ours 4.1225 26.596 138.21 857.60 2858.1
AF [2] 16.046 149.52 958.72 4321.8 $>$$9$
Enhanced AF 7.1205 53.508 327.96 1340.8 4929.3
Moss [29] 13.847 732.62 $>$$9$ $>$$9$ $>$$9$
Figure 10: The number of vertices for an arrangement in $\mathbb{R}^{4}$(LHS), $\mathbb{R}^{5}$(RHS) with 10 hyperplanes.
Both the Enhanced AF algorithm and ours improve Moss method, Avis and Fukuda’s method significantly. One highlight is that reversing the Zero rule is much better than reversing the Criss-Cross rule in the Enhanced AF.
Additionally, from each column in Table 7 and Figures 10, it can be observed that as the number of dimensions and vertices increases, the improvement becomes even more pronounced for any number of hyperplanes.
A three-layer ReLU neural network: We have built a neural network model with an input layer consisting of $7$ nodes to handle data from $\mathbb{R}^{3}$. Each node in the input layer receives a $3$-dimensional input. The hidden layer comprises $5$ nodes, each utilizes a weight vector with $7$ dimensions along with a bias term. These layers employ the Rectified Linear Unit (ReLU) activation function to introduce non-linearity into the network. In this network, all weight vectors and bias terms are randomly generated using the rand function in MATLAB. Firstly, we use the auxiliary linear programming to find the partition of the seven neurons in the input layer, resulting in a total of $64$ cells in $\mathbb{R}^{3}$. As the activation state of each cell is different, $5$ hyperplanes corresponding to the nodes in the hidden are also different. Next, we use algorithms to find out the total number of vertices of these $64$ different hyperplane arrangements with $12=7+5$ hyperplanes in $\mathbb{R}^{3}$. We randomly order these arrangements. The partial results are provided in Tables 8 and 9 and the overall result can be found in our link [^2] https://github.com/Github-DongZelin/Examples-and-Command-of-the-algorithm, $-1$ in the first arrangement means that none of the seven neurons in the input layer is active.
Method 1 4 7 11 14 17
0 Ours -1 185 185 220 220 185
AF [2]
Enhanced AF
@@ 0 Ours -1 185 185 220 220 185
20 Ours 220 220 220 220 220 110
AF [2]
Enhanced AF
@@ 20 Ours 220 220 220 220 220 110
40 Ours 220 220 220 220 220 220
AF [2]
Enhanced AF
@@ 40 Ours 220 220 220 220 220 220
Method 1 4 7 11 14 17
0 Ours -1 16.093 16.707 18.034 18.113 16.220
AF [2] -1 137.06 138.71 166.72 170.02 121.42
Enhanced AF -1 51.536 51.644 62.323 62.984 51.228
Moss [29] -1 146.06 148.13 218.69 223.03 149.95
20 Ours 18.054 18.158 18.151 18.052 17.956 10.935
AF [2] 161.31 171.20 173.15 166.21 175.76 70.709
Enhanced AF 62.794 62.485 63.56505 62.676 62.606 30.584
Moss [29] 220.469 223.52 227.02 227.72 227.44 52.524
40 Ours 24.996 24.178 18.661 18.641 18.723 18.651
AF [2] 198.09 237.25 170.78 179.18 179.57 178.85
Enhanced AF 65.223 84.506 64.889 64.750 65.086 65.466
Moss [29] 252.49 266.63 234.57 241.73 238.49 240.92
@@
From Table 8, it can be seen that the vertices in these hyperplane arrangements are highly degenerate. In a simple arrangement formed by 12 hyperplanes in $\mathbb{R}^{3}$, there should be $220$ vertices. However, several examples do not reach this value. Therefore, this series of examples can test the effectiveness of our method in highly degenerate hyperplane arrangement.
The following Figures 11 and Figures demonstrate the process of searching vertices in degenerated and non-degenerated arrangements, respectively, where we switch the order of color coverage in RHS of Figure 11 for better visualization. It can be seen that, in all cases, ours and the Enhanced AF are the most and the second most efficient, respectively.
Figure 11: Vertices in 17-th(LHS) and 37-th(RHS) arrangement
Figure 12: Vertices in 21-st(LHS) and 41-st(RHS) arrangement
Result Analysis: As Tables 3-9 shows, in arrangements with a smaller amount of hyperplanes, both ours and the Moss method demonstrate higher efficiency, with the Moss method even achieving the highest efficiency in some cases. However, as the sample size $n$ and the dimensionality $d$ increase, the efficiency of the Moss method declines significantly, while ours consistently maintains high efficiency. Additionally, when comparing the performance of the Enhanced AF and the original AF, which typically rank second and third, it is clear that the Enhanced AF is superior to the original AF. This finding corroborates the advantages of directly employing the if-and-only-if condition to determine valid reverse Criss-Cross pivots in Appendix.
Moreover, as Figures 10-12 show, for each arrangement, in the beginning, both ours and the Moss method exhibit higher rates in finding vertices than the rest two algorithms. As the search continues, the rates of our algorithm and the Enhanced AF remain relatively stable, the rate of the original AF fluctuates, and the rate of the Moss method gradually declines. Due to the stability, the Enhanced AF and ours enjoy the high computational efficiency than the original AF. The gradual decline in the search rate reveals that the Moss method is handicapped in handling complex cases.
The reason for the stability of the Enhanced AF and ours lies in the if-and-only-if conditions for detecting valid reverse pivots, where the computational cost at each position of any dictionary is the same. In contrast, the variability in efficiency of the original AF arises from the varying computational costs at each position, leading to a significant amount of unnecessary calculations. The decreased efficiency of the Moss can be attributed to the increased time required for checking duplicates, as the number of discovered dictionaries grows. Consequently, it performs well with a small sample size but poorly with a larger sample size.
7 Conclusion
In this paper, we present the Zero rule for the VE problem of hyperplane arrangements. Compared to the classical Criss-Cross rule, the Zero rule enjoys several computationally friendly properties: i) It can get rid of the objective function, which saves the computation. ii) The condition between the Zero rule and its valid reverse is sufficient and necessary, which is more efficient when determining valid reverse pivots. iii) In $\mathbb{R}^{d}$, the number of pivot steps is at most $d$, which can transform the unpredictable depth into a certain maximum depth, further saving the computation. iv) Its terminal within the entire arrangement is unique, which is convenient for algorithmic design. Because of these properties, the complexity of the VE algorithm using the Zero rule is $\mathcal{O}(n^{2}d^{2}(v-v_{d})+ndv_{d})$. Moreover, it could be as low as $\mathcal{O}(nd^{4}v)$ for simple arrangements, which greatly improves the state-of-the-art algorithms. Systematic and comprehensive experiments have confirmed not only the efficiency of using the if-and-only-if-condition to determine valid reverse pivot but also the efficiency of the VE algorithm using the Zero rule. In the future, more efforts should be made to further escalate the algorithmic efficiency.
8 Acknowledgement
Appendix. Analysis about the checking of valid reverse pivot
To enhance the efficiency of the algorithm by Avis and Fukuda [2], we provide the necessary and sufficient conditions for determining whether the pivot $(s,r)$ on a dictionary is a valid reverse Criss-Cross pivot, as illustrated in Figure 13. These conditions represent the minimal requirements to confirm a valid reverse Criss-Cross pivot.
Proposition 8.
Let $x_{B}=\bar{A}x_{N}$ be an arbitrary dictionary with the objective function. Then pivot $(s,r)$, where $s\in B_{\neq f}$ and $r\in N_{\neq g}$, is a valid reverse Criss-Cross pivot if and only if either (a)